# Decay of solutions of nonlinear Dirac equations

### v2

@@ 1.4. Solitary waves
Theorem 1.4 addresses the problem of decay in a subspace of the partial wave subspace where small solitary waves are not present. To complement the previous literature we discuss the case of coherent structures (see [64, 34]).
The existence and stability of solitary waves are problems that have been widely studied, however, it is not completely well understood. The stability problem, in particular, has been extensively explored in both one-dimensional (1D) and three-dimensional (3D) cases. For the 1D case, notable results have been obtained concerning the asymptotic stability of solitary waves in certain Dirac-type models, often under specific restrictions on the types of perturbations considered.
@@
Regarding asymptotic stability, Comech, Van Phan, and Stefanov [24] addressed the 1D nonlinear Dirac equation with scalar self-interaction involving quintic and higher-order nonlinearities, known as the Gross-Neveu model. They established the asymptotic stability of solitary waves for a range of frequencies. In their analysis, they considered “even” perturbations, excluding in this way translations and eigenvalues of the form ±$\pm 2wi$.
Furthermore on the 1D Dirac equation in (1.8). In [23], described the existence and explicit form of gap solitary waves solutions but also studied the spectral stability of such solutions. Pelinovsky and Shimabukuro [56], for the Massive Thirring Model, proved that the solitary waves are orbitally stable on $H^{1}$. After, the same authors and Contreras [25], proved that the solitary waves are orbitally stable on $L^{2}$. The authors rely on the integrability of the Massive Thirring Model, thanks to the inverse scattering transform they find a new conserved quantity and are able to use the auto-Bäcklund transformation. Regarding the asymptotic stability, in [57] the authors proved dispersive decay estimates and used them to prove asymptotic stability of small gap solitons with quintic and higher-order nonlinear terms. For the 3D Dirac equation, the standing wave had been described using a symmetry decomposition (see [14, 63, 6, 51]), and the stability properties have been studied to these solitary waves.
Cazenave and Vazquez [20] proved the existence of stationary states for the 3D nonlinear Dirac equation with a nonlinearity $-g(\overline{\psi}\psi)$. Following the Wakano ansatz [67] and Soler [62], they obtained solutions that are separable in spherical coordinates and exponentially decaying. This result was generalized for a wide variety of nonlinearities by Merle [51], e.g., nonlinearities that are not necessarily increasing or bounded below their value in zero.
@@
Recalling the connection between NLS and nonlinear Dirac equation. We shall highlight that the Soler model shares some interesting but complex properties with the focusing cubic NLS. In the focusing cubic NLS, it has been proved that there is a family of multi-solitons which are arbitrarily small perturbations in $H^{1}$, of the solitary wave $\sqrt{2}\operatorname{sech}(x)$ (see [47] and the reference therein). These features have been described by Boussaïd and Comech in [12] for the 3D Soler model. Furthermore, they showed that the bi-frequency solitary waves are associated with the Bogoliubov ${\bf SU}(N/2,N/2)$ symmetry, giving some lines about how it is the proper way to understand the asymptotic stability of these solutions. In [13], the same authors and Kulkarni proved that bi-frequency solitary waves might also possess linear stability properties similar to those present in the one-frequency solitary wave.
@@ 1.7. Examples
${\bf W}:=\frac{1}{2}G(X^{T}AX),$ (1.28)
with $G:\mathbb{R}\to\mathbb{R}$ a given polynomial nonlinearity of order at least 2 and satisfying $G(0)=0$. Define $G^{\prime}=:g$, so that $|g(s)|\lesssim s^{p}$, with $p\geq 1$, $p\in{\mathbb{N}}$. Then, we obtain
$W_{jk}=g\left(\sum_{l,m=1,2}a_{lm}\phi_{lm}^{2}\right)a_{jk}\phi_{jk},\quad\mbox{ for }j,k=1,2.$
Then (1.20) is satisfied. Therefore, the previous examples provides an example of a class of nonlinearities for which Theorem 1.4 holds. Notice that Soler’s nonlinearity (1.21) does not fit into (1.28).
@@ 3.1. Identities for the $n$ dimensional Dirac equation
[…]
@@ 4.1. First computations
This means that the new functional $\mathcal{J}^{j}_{t_{0}}(t)$ is decreasing in $[2,t_{0}]$. Therefore, we have
[…]
@@
00 0 $\leq\int\varphi_{j}\left(\dfrac{x+\boldsymbol{\theta}t_{0}-\boldsymbol{\tilde{\theta}}(t_{0}-t)}{\lambda}\right)\left[\boldsymbol{u}_{1}\boldsymbol{u}_{1}+\boldsymbol{u}_{2}\boldsymbol{u}_{2}\right](t_{0},x)$
@@
[…]
@@
[…]
@@
[…]
since $\frac{1}{\lambda}$ does not integrable in $[10,\infty)$, this concludes the proof of this corollary. ∎
@@
It is interesting to discuss the meaning of Lemma 6.1 in the integrable case $W=|u|^{2}|v|^{2}$ (1.8)-(1.27). Since $\partial_{a}W_{1}-\partial_{c}W_{2}=\frac{1}{2}(|\psi_{1}|^{2}-|\psi_{2}|^{2})\neq 0$ in principle, the term $-\left(\partial_{c}W_{2}-\partial_{a}W_{1}\right)\partial_{x}\psi_{2}$ in (6.2) survives. This term is doubly dangerous, since first: it considers the opposite variable $\psi_{2}$, and second, it considers the derivative of that variable. Naturally, the huge soliton solution manifold represented by (1.7) is deeply related to the existence of this additional term, and it is not present in classical NLKG models. Consequently, when asking the Cauchy-Riemann condition (1.12), we are discarding the integrable and all other cases of nonlinearities where mixed spatial partial derivatives appear, which are suitable conditions for the emergence of Dirac’s type solitary waves. We also recall that (1.12) is part of a sufficient condition, not enough to conclude (since one also needs a sort of conservation law in the problem), but it is also necessary, in the sense that the integrable case is a counterexample under the absence of (1.12).
@@ 6.1. Virial identities I
From (6.15) there exists a sequence $t_{n}\to+\infty$ such that $\mathcal{H}(t_{n})\to 0$. Let $t\in\mathbb{R}$. Integrating (6.16) on $[t,t_{n}]$ and passing to the limit as $n\to+\infty$ we obtain
[…]
@@ 7.1. Virial Identities II
First, we consider the term $A$ defined in (7.10). Similar to the 1D case, this time the term $A$ has a richer structure and under the assumption (1.20) we recover a full derivative. Using that $W_{i1}=\Re\partial_{\overline{\phi_{i}}}{\bf W}$ and $W_{i2}=\Im\partial_{\overline{\phi_{i}}}{\bf W}$, one obtains
$A=$ $~{}{}\frac{1}{2}\int\varphi\left(\overline{\partial_{r}\phi}_{1}\partial_{\overline{\phi}_{1}}{\bf W}+\partial_{r}\phi_{1}\overline{\partial_{\overline{\phi}_{1}}{\bf W}}+\overline{\partial_{r}\phi}_{2}\partial_{\overline{\phi}_{2}}{\bf W}+\partial_{r}\phi_{2}\overline{\partial_{\overline{\phi}_{2}}{\bf W}}\right)$
$+\int\varphi^{\prime}\big{(}W_{11}\phi_{11}+W_{12}\phi_{12}+W_{21}\phi_{21}+W_{22}\phi_{22}\big{)}$
$=$ $~{}{}\frac{1}{2}\int\varphi\left(\overline{\partial_{r}\phi}_{1}\partial_{\overline{\phi}_{1}}{\bf W}+\partial_{r}\phi_{1}\partial_{\phi_{1}}\overline{{\bf W}}+\overline{\partial_{r}\phi}_{2}\partial_{\overline{\phi}_{2}}{\bf W}+\partial_{r}\phi_{2}\partial_{\phi_{2}}\overline{{\bf W}}\right)$
$+\int\varphi^{\prime}\big{(}W_{11}\phi_{11}+W_{12}\phi_{12}+W_{21}\phi_{21}+W_{22}\phi_{22}\big{)}$
$=$ $~{}{}\frac{1}{2}\int\varphi\partial_{r}\left({\bf W}\right)+\int\varphi^{\prime}\big{(}W_{11}\phi_{11}+W_{12}\phi_{12}+W_{21}\phi_{21}+W_{22}\phi_{22}\big{)},$
since ${\bf W}$ is real-valued. Finally, recalling that $|{\bf W}|\leq C|\phi|^{p}$ for $p\geq 4$, we obtain
$|mA|\lesssim\int\varphi^{\prime}|\phi|^{p}.$ (7.12)
Now, we will focus on the terms associated to the nonlinear terms presents on Corollary 7.4, $B$ in (7.11). Let us consider the following decomposition of $B$,
$B=$ $~{}{}2\int\varphi\big{(}\partial_{r}W_{11}\partial_{r}\phi_{21}+\partial_{r}W_{12}\partial_{r}\phi_{22}+\partial_{r}W_{21}\partial_{r}\phi_{11}+\partial_{r}W_{22}\partial_{r}\phi_{12}\big{)}$
@@
$\varphi^{\prime}=\frac{\sqrt{r}(r+3)}{2(1+r)^{2}},$
@@
Similarly, recalling (7.12), one gets
$|mA|\lesssim$ $~{}{}\delta^{2}\int\frac{\sqrt{r}(r+3)}{2(1+r)^{4}}|\phi|^{2}\lesssim\delta^{2}\int\frac{r^{2}(r+3)}{2r^{3/2}(1+r)^{4}}|\phi|^{2}.$
Notice that for $B_{1}$, since $\varphi\geq 0$, we have
@@
$\leq$ $~{}{}\varepsilon\int\frac{\sqrt{r}}{(1+r)^{2}}|\nabla\phi|^{2}+C_{\varepsilon}\int\sqrt{r}|W|^{2},$
where $W=(W_{1},W_{2})$.
@@
$~{}{}-C\int\sqrt{r}|W|^{2}+B_{2}+B_{3}.$
Now, since $\sup_{t\in\mathbb{R}}\|\phi(t)\|_{L^{\infty}}\lesssim\delta$, using the estimate $|\phi|\lesssim\frac{\delta}{(1+r)}$ [49] we get
$-C\int\sqrt{r}|W|^{2}\gtrsim-\int\frac{r^{2}}{r^{3/2}}|\phi|^{6}\gtrsim-\delta^{4}\int\frac{r^{2}|\phi|^{2}}{r^{3/2}(1+r)^{4}}.$
Consequently, (7.13) becomes
$\frac{d}{dt}\mathcal{K}\geq$ $~{}{}\frac{3}{4}\int\frac{\sqrt{r}}{2(1+r)}|\nabla\phi|^{2}+\int\frac{(r^{3}+r^{2}+r+1)}{64r^{3/2}(1+r)^{4}}|\phi|^{2}+B_{2}+B_{3}.$
Finally, for $B_{2}$ and $B_{3}$, we notice that
@@
$-\frac{1}{2}\bigg{(}\varphi^{\prime\prime}-2\frac{\varphi^{\prime}}{r}\bigg{)}=\frac{5r^{2}+22r+9}{8\sqrt{r}(1+r)^{3}},$
and similar to $B_{1}$, we obtain
@@
From this estimate and the boundedness of $\mathcal{K}$ for all times we obtain the integral bound
[…]
@@
Now we consider the regions $10<r<$ 1 and $r\geq 1$ separately. First of all,
@@ 7.2. End of proof
Recall that (7.14) implies that there exists a sequence $t_{n}\to\infty$ such that $\mathcal{H}(t_{n})\to 0$. Integrating the inequality above on $[t,t_{n}]$ we see that
[…]
@@ Appendix A Computations of virial identities
×$~{}{}\quad\times\big{(}-\partial_{x}^{2}\phi_{12}-m\partial_{x}\phi_{22}+\partial_{x}(W_{22})+m\partial_{x}\phi_{22}+m^{2}\phi_{12}-mW_{12}\big{)}$
@@
×$~{}{}\quad\times(\partial_{x}\phi_{21}+m\phi_{11}).$
@@
×$\qquad\times(\partial_{x}\phi_{11}+m\phi_{21})$
@@
×$\qquad\times\big{(}\partial_{x}^{2}\phi_{22}+m\partial_{x}\phi_{12}-\partial_{x}(W_{12})-m\partial_{x}\phi_{12}-m^{2}\phi_{22}+mW_{22}\big{)}$
@@
×$\qquad\times(\partial_{x}\phi_{11}+m\phi_{21})$
@@
×$~{}{}\quad\times\big{(}-\partial_{x}^{2}\phi_{21}-m\partial_{x}\phi_{11}+\partial_{x}(W_{11})+m\partial_{x}\phi_{11}+m^{2}\phi_{21}-mW_{21}\big{)}$
@@
×$~{}{}\quad\times(\partial_{x}\phi_{12}+m\phi_{22}).$
@@
×$\qquad\times\left(\left(\partial_{r}+\frac{2}{r}\right)\phi_{21}+m\phi_{11}\right)$
@@
×$\qquad\times\left(\left(\partial_{r}+\frac{2}{r}\right)(-\partial_{r}\phi_{12}-m\phi_{22}+W_{22})+m\left(\left(\partial_{r}+\frac{2}{r}\right)\phi_{22}+m\phi_{12}-W_{12}\right)\right).$
@@
×$~{}{}\qquad\times(\partial_{r}\phi_{12}+m\phi_{22})$
@@
×$~{}{}\qquad\times\left(\partial_{r}\left(\left(\partial_{r}+\frac{2}{r}\right)\phi_{21}+m\phi_{11}-W_{11}\right)-m(\partial_{r}\phi_{11}+m\phi_{21}-W_{21})\right).$

### v4

@@ 1.4. Solitary waves
Theorem 1.4 addresses the problem of decay in a subspace of the partial wave subspace where small solitary waves are not present. To complement the previous literature, we discuss the case of coherent structures (see [64, 34]).
The existence and stability of solitary waves are problems that have been widely studied; however, it is not completely well understood. The stability problem, in particular, has been extensively explored in both one-dimensional (1D) and three-dimensional (3D) cases. For the 1D case, notable results have been obtained concerning the asymptotic stability of solitary waves in certain Dirac-type models, often under specific restrictions on the types of perturbations considered.
@@
Regarding asymptotic stability, Comech, Van Phan, and Stefanov [24] addressed the 1D nonlinear Dirac equation with scalar self-interaction involving quintic and higher-order nonlinearities, known as the Gross-Neveu model. They established the asymptotic stability of solitary waves for a range of frequencies. In their analysis, they considered “even” perturbations, excluding in this way translations and eigenvalues of the form $\pm 2wi$.
Furthermore, on the 1D Dirac equation in (1.8). In [23], described the existence and explicit form of gap solitary waves solutions but also studied the spectral stability of such solutions. Pelinovsky and Shimabukuro [56], for the Massive Thirring Model, proved that the solitary waves are orbitally stable on $H^{1}$. After, the same authors and Contreras [25], proved that the solitary waves are orbitally stable on $L^{2}$. The authors rely on the integrability of the Massive Thirring Model, thanks to the inverse scattering transform they find a new conserved quantity and are able to use the auto-Bäcklund transformation. Regarding the asymptotic stability, in [57] the authors proved dispersive decay estimates and used them to prove asymptotic stability of small gap solitons with quintic and higher-order nonlinear terms. For the 3D Dirac equation, the standing wave had been described using a symmetry decomposition (see [14, 63, 6, 51]), and the stability properties have been studied these solitary waves.
Cazenave and Vazquez [20] proved the existence of stationary states for the 3D nonlinear Dirac equation with a nonlinearity $g(\overline{\psi}\psi)$. Following the Wakano ansatz [67] and Soler [62], they obtained solutions that are separable in spherical coordinates and exponentially decaying. This result was generalized for a wide variety of nonlinearities by Merle [51], e.g., nonlinearities that are not necessarily increasing or bounded below their value in zero.
@@
Recalling the connection between NLS and the nonlinear Dirac equation. We shall highlight that the Soler model shares some interesting but complex properties with the focusing cubic NLS. In the focusing cubic NLS, it has been proved that there is a family of multi-solitons which are arbitrarily small perturbations in $H^{1}$, of the solitary wave $\sqrt{2}\operatorname{sech}(x)$ (see [47] and the reference therein). These features have been described by Boussaïd and Comech in [12] for the 3D Soler model. Furthermore, they showed that the bi-frequency solitary waves are associated with the Bogoliubov ${\bf SU}(N/2,N/2)$ symmetry, giving some lines about how it is the proper way to understand the asymptotic stability of these solutions. In [13], the same authors and Kulkarni proved that bi-frequency solitary waves might also possess linear stability properties similar to those present in the one-frequency solitary wave.
@@ 1.7. Examples
$(W_{1},W_{2}):=g(X^{T}AX)X,$ (1.28)
with $g:\mathbb{R}\to\mathbb{R}$ a given polynomial nonlinearity of order at least 1, satisfying $g(0)=0$. Thus, it holds that $|g(s)|\lesssim s^{p}$, with $\geq 1, pp\in{\mathbb{N}}$. Then, we obtain
$W_{jk}=g\left(\sum_{l,m=1,2}a_{lm}\phi_{lm}^{2}\right)\phi_{jk},\quad\mbox{ for }j,k=1,2.$
This type of nonlinearity satisfies (1.20). Therefore, this construction provides a class of nonlinearities for which Theorem 1.4 holds. Notice that Soler’s nonlinearity (1.22) fits into (1.28), and therefore satisfies the hypotheses of Theorem 1.4.
@@ 3.1. Identities for the $n$ dimensional Dirac equation
$\mathcal{I}=\frac{1}{\mu(t)}\int\varphi\left(\frac{x+\rho(t)}{\lambda(t)}\right)\psi^{\dagger}(t,x)\psi(t,x)dx.$
@@ 4.1. First computations
This means that the new functional $\mathcal{J}^{j}_{t_{0}}(t)$ is decreasing in $2,t_{0}]$. Therefore, we have
$\int_{2}^{t_{0}}\dfrac{d}{dt}\mathcal{J}^{j}_{t_{0}}(t)dt=\mathcal{J}^{j}_{t_{0}}(t_{0})-\mathcal{J}^{j}_{t_{0}}(2)\leq 0\implies\ \mathcal{J}^{j}_{t_{0}}(t_{0})\leq\mathcal{J}^{j}_{t_{0}}(2).$
@@
$0$ $\leq\int\varphi_{j}\left(\dfrac{x+\boldsymbol{\theta}t_{0}-\boldsymbol{\tilde{\theta}}(t_{0}-t)}{\lambda}\right)\left[\boldsymbol{u}_{1}\boldsymbol{u}_{1}+\boldsymbol{u}_{2}\boldsymbol{u}_{2}\right](t_{0},x)$
@@
$\limsup_{t\to\infty}\int\varphi_{j}\left(\dfrac{x+\boldsymbol{\theta}t_{0}-\boldsymbol{\tilde{\theta}}(t_{0}-t)}{L}\right)\left[\boldsymbol{u}_{1}\boldsymbol{u}_{1}+\boldsymbol{u}_{2}\boldsymbol{u}_{2}\right](t,x)dx=0.$
@@
    $\int_{10}^{\infty}\frac{1}{\lambda(t)}\int_{\mathbb{R}}\operatorname{sech}^{2}\left(\frac{x}{\lambda(t)}\right)(|u|^{2}+|v|^{2})(t,x)dxdt\lesssim 1.$
@@
$\int_{10}^{\infty}\frac{1}{\lambda}\int\phi^{\prime}\left(\frac{x}{\lambda}\right)(|u|^{2}+|v|^{2})dxdt<\infty,$
since $\frac{1}{\lambda}$ does not integrable in $10,\infty)$, this concludes the proof of this corollary. ∎
@@
It is interesting to discuss the meaning of Lemma 6.1 in the integrable case $W=|u|^{2}|v|^{2}$ (1.8)-(1.27). Since $\partial_{a}W_{1}-\partial_{c}W_{2}=\frac{1}{2}(|\psi_{1}|^{2}-|\psi_{2}|^{2})\neq 0$ in principle, the term $\left(\partial_{c}W_{2}-\partial_{a}W_{1}\right)\partial_{x}\psi_{2}$ in (6.2) survives. This term is doubly dangerous, since first: it considers the opposite variable $\psi_{2}$, and second, it considers the derivative of that variable. Naturally, the huge soliton solution manifold represented by (1.7) is deeply related to the existence of this additional term, and it is not present in classical NLKG models. Consequently, when asking the Cauchy-Riemann condition (1.12), we are discarding the integrable and all other cases of nonlinearities where mixed spatial partial derivatives appear, which are suitable conditions for the emergence of Dirac’s type solitary waves. We also recall that (1.12) is part of a sufficient condition, not enough to conclude (since one also needs a sort of conservation law in the problem), but it is also necessary, in the sense that the integrable case is a counterexample under the absence of (1.12).
@@ 6.1. Virial identities I
From (6.15) there exists a sequence $t_{n}\to+\infty$ such that $\mathcal{H}(t_{n})\to 0$. Let $t\in\mathbb{R}$. Integrating (6.16) on $t,t_{n}]$ and passing to the limit as $n\to+\infty$ we obtain
$\mathcal{H}(t)\lesssim\int_{t}^{+\infty}\int\operatorname{sech}\left(x\right)\left[|\nabla\phi|^{2}+|\phi|^{2}\right]dt.$
@@ 7.1. Virial Identities II
First, we consider the term $A$ defined in (7.10). Unlike the 1D case, the term $A$ has a more general structure. However, the assumption (1.20), along with Lemma 7.5, is enough to deal with it. Indeed, using the fact that the nonlinearity satisfies (1.20), one obtains
$|A|\lesssim$ $~{}\int\varphi|\phi|^{3}|\nabla\phi|+\int|\varphi^{\prime}||\phi|^{4}$
$\lesssim$ $~{}\int\varphi(|\nabla\phi|^{2}|\phi|+|\phi|^{5})+\int|\varphi^{\prime}||\phi|^{4}$
$\lesssim$ $~{}\delta\int\frac{\varphi}{1+r}|\nabla\phi|^{2}++\delta^{3}\int\frac{\varphi}{(1+r)^{3}}|\phi|^{2}+\delta^{2}\int\frac{|\varphi^{\prime}|}{(1+r)^{2}}|\phi|^{2},$
since $(W_{1},W_{2})$ is a pure power nonlinearity. Finally, rearranging the terms, we obtain
$|mA|\lesssim\delta\int\frac{\varphi}{1+r}|\nabla\phi|^{2}+\delta^{2}\int\bigg(\frac{\varphi}{(1+r)^{3}}+\frac{|\varphi^{\prime}|}{(1+r)^{2}}\bigg)|\phi|^{2}.$ (7.12)
Now, we will focus on the terms associated with the nonlinear terms present in Corollary 7.4, $B$ in (7.11). Let us consider the following decomposition of $B$,
@@
$\varphi^{\prime}=\frac{\sqrt{r}(r+3)}{2(1+r)^{2}},$ (7.15)
@@
Similarly, recalling (7.12) along with (7.14) and (7.15), one gets
$|mA|\lesssim$ $~{}\delta\int\frac{\varphi}{1+r}|\nabla\phi|^{2}+\delta^{2}\int\bigg(\frac{\varphi}{(1+r)^{3}}+\frac{|\varphi^{\prime}|}{(1+r)^{2}}\bigg)|\phi|^{2}$
$\lesssim$ $~{}\delta\int\frac{r^{3/2}}{(1+r)^{2}}|\nabla\phi|^{2}+\delta^{2}\int\bigg(\frac{r^{3/2}}{(1+r)^{4}}+\frac{\sqrt{r}(r+3)}{2(1+r)^{4}}\bigg)|\phi|^{2}$
$\lesssim$ $~{}\delta\int\frac{r^{3/2}}{(1+r)^{2}}|\nabla\phi|^{2}+\delta^{2}\int\frac{3\sqrt{r}(r+1)}{2(1+r)^{4}}|\phi|^{2}.$
Notice that for $B_{1}$, since $\varphi\geq 0$ and $W=(W_{1},W_{2})$ satisfies (1.20), we have
@@
$\lesssim$ $~{}\int\frac{\varphi}{r}|W||\nabla\phi|$
$\lesssim$ $~{}\int\frac{\varphi}{r}(|\nabla\phi|^{2}|\phi|+|\phi|^{2p-1})$
$\lesssim$ $~{}\delta\int\frac{\sqrt{r}}{(1+r)^{2}}|\nabla\phi|^{2}+\delta^{3}\int\frac{\sqrt{r}}{(1+r)^{4}}|\phi|^{2}.$
@@
$~{}+B_{2}+B_{3}.$
Finally, let us focus on $B_{2}$ and $B_{3}$ in (7.13). By using the estimate $|\phi|\lesssim\frac{\delta}{(1+r)}$ [49] and the fact that $\sup_{t\in\mathbb{R}}\|\phi(t)\|_{L^{\infty}}\lesssim\delta$, we obtain
@@
$-\frac{1}{2}\bigg(\varphi^{\prime\prime}-2\frac{\varphi^{\prime}}{r}\bigg)=\frac{5r^{2}+22r+9}{8\sqrt{r}(1+r)^{3}}.$
Similarly to $B_{1}$, we obtain
@@
From this estimate and the boundedness of $\mathcal{K}$ for all times, we obtain the integral bound
$\int_{0}^{\infty}\int\left(\dfrac{\sqrt{r}|\nabla\phi|^{2}}{(1+r)}+\dfrac{|\phi|^{2}}{r^{3/2}(1+r)}\right)dr\,dt\lesssim 1.$ (7.17)
@@
Now we consider the regions $0<r<1$ and $r\geq 1$ separately. First of all,
@@ 7.2. End of proof
Recall that (7.17) implies that there exists a sequence $t_{n}\to\infty$ such that $\mathcal{H}(t_{n})\to 0$. Integrating the inequality above on $t,t_{n}]$ we see that
$|\mathcal{H}(t_{n})-\mathcal{H}(t)|=$ $\left|\int_{t}^{t_{n}}\dfrac{d}{dt}\mathcal{H}(s)ds\right|\leq\int_{t}^{t_{n}}\left|\dfrac{d}{dt}\mathcal{H}(s)ds\right|\lesssim\int_{t}^{t_{n}}\mathcal{H}(s)ds,$
@@ Appendix A Computations of virial identities
$~{}\quad\times\big(-\partial_{x}^{2}\phi_{12}-m\partial_{x}\phi_{22}+\partial_{x}(W_{22})+m\partial_{x}\phi_{22}+m^{2}\phi_{12}-mW_{12}\big)$
@@
$~{}\quad\times(\partial_{x}\phi_{21}+m\phi_{11}).$
@@
$\qquad\times(\partial_{x}\phi_{11}+m\phi_{21})$
@@
$\qquad\times\big(\partial_{x}^{2}\phi_{22}+m\partial_{x}\phi_{12}-\partial_{x}(W_{12})-m\partial_{x}\phi_{12}-m^{2}\phi_{22}+mW_{22}\big)$
@@
$\qquad\times(\partial_{x}\phi_{11}+m\phi_{21})$
@@
$~{}\quad\times\big(-\partial_{x}^{2}\phi_{21}-m\partial_{x}\phi_{11}+\partial_{x}(W_{11})+m\partial_{x}\phi_{11}+m^{2}\phi_{21}-mW_{21}\big)$
@@
$~{}\quad\times(\partial_{x}\phi_{12}+m\phi_{22}).$
@@
$\qquad\times\left(\left(\partial_{r}+\frac{2}{r}\right)\phi_{21}+m\phi_{11}\right)$
@@
$\qquad\times\left(\left(\partial_{r}+\frac{2}{r}\right)(-\partial_{r}\phi_{12}-m\phi_{22}+W_{22})+m\left(\left(\partial_{r}+\frac{2}{r}\right)\phi_{22}+m\phi_{12}-W_{12}\right)\right).$
@@
$~{}\qquad\times(\partial_{r}\phi_{12}+m\phi_{22})$
@@
$~{}\qquad\times\left(\partial_{r}\left(\left(\partial_{r}+\frac{2}{r}\right)\phi_{21}+m\phi_{11}-W_{11}\right)-m(\partial_{r}\phi_{11}+m\phi_{21}-W_{21})\right).$